OpenStudy (coolshubhs):

If a,b,c in A.P. (progression) prove that : (c-a)^2 = 4(a-b)(b-a)

OpenStudy (anonymous):

LET a,b,c be a-d,a,a+d put these values for a,b,c in above expression and show L.H.S=R.H.S

OpenStudy (klimenkov):

мау ве тне 5есоиd bгаскет !и R.H.S = (с-b)

OpenStudy (coolshubhs):

I think we have to use the formula : b-a=c-b

OpenStudy (anonymous):

a-c=0 and a-b=0 a=b=c

OpenStudy (anonymous):

A^2+B^2=0 -> A=0 and B=0

OpenStudy (anonymous):

(a-c)2+4(a-b)2=0

OpenStudy (klimenkov):

@mahmit2012 why a-c=0 and a-b=0 in A.P.?

OpenStudy (anonymous):

summation of two non negative terms lways is non negative. So if it is equal to zero, all term should be zero.

OpenStudy (coolshubhs):

but we cant say a-c=0 :/

OpenStudy (anonymous):

a b c = 2 4 6 -> 4^2 not equals 4 (-2)(2)

OpenStudy (anonymous):

a, a+k, a+2k same problem

OpenStudy (coolshubhs):

it's a question from arithmetic progression.. we can use the formula b-a=c-b to prove ...

OpenStudy (anonymous):

Something wrong maybe it is 4(a-b)(a-b) or something...

OpenStudy (coolshubhs):

the previous question was .. (c-a)^2=4(b^2-ac) I have done that.

OpenStudy (anonymous):

Yes, that one seems to work OK